A fruit drink concentrate is 15% water. How many liters of pure water should be added to 12 L of concentrate to produce a mixture that is 50% water?
Answer by richwmiller(17219)

.15*12+1*x=.5(12+x)
1.8+x=6+.5x
18+10x=60+5x
10x-5x=60-18
5x=42
x=42/5
x=8.4 liters of pure water
check
1.8+8.4=.5(20.4)
10.2=10.2
ok